a man standing in the middle of a pedestrian bridge over a highway sights two stranded motorists, one on each side of the bridge. From his position, the car on one side is at an angle of depression of 15 and the car on the other side is at an angle of depression of 28. the bridge is 20 feet high. How far apart are the two cars?
Answer by Boreal(15235)

This must be drawn.
tan 15=20/x
x=20/tan 15=74.64 feet
similarly y= 20/ tan 28
y=37.61 feet
The cars are 112.25 feet apart.